# Verifiable type-III seesaw and dark matter in a gauged $\boldsymbol{U(1)_{\rm B-L}}$ symmetric model

### v2

@@ I Introduction
The Standard Model (SM) of particle physics, despite its remarkable success, leaves several fundamental questions unanswered. Two of the most pressing mysteries are the nature of dark matter (DM) and the origin of neutrino masses. Observational evidence from astrophysics and cosmology, including galaxy rotation curves, gravitational lensing, and cosmic microwave background data, strongly suggests that DM constitutes approximately $85\%$ of the Universe’s matter content and $26.8\%$ of its energy density [1, 2]. However, the SM lacks a suitable candidate to explain the DM of the universe.
@@
Concurrently, the discovery of neutrino oscillations [3, 4, 5, 6, 7] has conclusively demonstrated that neutrinos possess non-zero masses, contradicting the SM prediction of massless neutrinos. This observation necessitates an extension of the SM to accommodate neutrino mass generation mechanisms. Various seesaw mechanisms have been proposed to explain the small neutrino masses, including the type I [8, 9, 10, 11, 12], type II [13, 14, 15, 16, 17], and type III [18] seesaw models.
@@
The potential connection between these two fundamental puzzles has sparked considerable interest in the particle physics community. Several beyond Standard Model (BSM) theories have been developed to address both issues within a unified framework. Among these, models based on the gauged $U(1)_{\rm B-L}$ symmetry have garnered significant attention due to their simplicity and potential for experimental verifiability. See for instance  [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29] and the references there in.
@@
In this paper, we present a comprehensive study of a $U(1)_{\rm B-L}$ gauged type-III seesaw model that simultaneously addresses the neutrino mass puzzle and provides a DM candidate. We explore the model’s theoretical foundations, analyze its phenomenological consequences, and discuss its implications for current and future experimental searches. The type-III seesaw mechanism [18], which introduces $SU(2)_{L}$ fermion triplets, offers a unique approach to neutrino mass generation. However, when embedded within a gauged $U(1)_{\rm B-L}$ symmetric setup, it introduces new anomalies that require additional particle content for cancellation. Intriguingly, the fermions needed for anomaly cancellation in this setup can serve as viable DM candidates, establishing a natural link between neutrino physics and DM. Specifically, after a careful assignment of $\rm B-L$ charges to the triplet fermions which successfully cancels the $SU(2)_{L}^{2}\times U(1)_{\rm B-L}$ anomalies, two additional chiral fermions are required to cancel the remaining $U(1)_{\rm B-L}^{3}$ and $(grav.)^{2}\times U(1)_{\rm B-L}$ anomalies. These fermions can combine to form a Dirac particle, which gets stabilized by the remnant $Z_{2}$ symmetry after the $U(1)_{\rm B-L}$ breaking and thus becomes a natural DM candidate.
@@
Here it is worth noting that, as the triplet fermions carry electroweak charge their production and decay signatures can be prominent at colliders as compared to the type-I seesaw scenario, and by the virtue of their $SU(2)_{L}$ interactions, they can be pair‑produced via electroweak processes with cross sections set by the triplet mass [30, 31]. In contrast, heavy right-handed neutrinos in type-I seesaw are SM singlets and are only produced via tiny mixing angles, making them essentially invisible at colliders. However, a $\Sigma$ can be copiously produced at the LHC/HL-LHC once kinematically allowed. The charged component $\Sigma^{\pm}$ decays to the neutral $\Sigma^{0}$ plus soft pions (or leptons) with a potentially long lifetime because of the phase space suppression due to the small mass difference arising because of electroweak corrections [32]. If $\Sigma^{\pm}$ is moderately long-lived, it yields a disappearing track signature in the detector. This striking signal, absent in Type-I models, can be searched for at the LHC/HL-LHC. Type-III models also readily produce displaced vertex events, and proposed detectors like MATHUSLA or future lepton–hadron colliders (LHeC, FCC-he) could detect such displaced decays [30, 31]. Thus, type-III seesaw implementations can be far more testable and constrained than the minimal Type-I seesaw.
@@
In addition, the model also predicts unique cosmological signatures in the form of stochastic gravitational waves (GWs). These GWs originate from first-order phase transitions during the breaking of the $U(1)_{\rm B-L}$ symmetry, a process central to the framework’s phenomenology. The resulting GW spectrum can be probed by current and future observatories such as LISA [33], DECIGO [34], and the Einstein Telescope [35]. These experiments provide a complementary avenue to test the model, as the GW signals are sensitive to the dynamics of the phase transition and the underlying particle physics parameters.
@@
The rest of this paper is organized as follows: In Section II, we provide a detailed description of our model, including the particle content, gauge symmetries, and the Lagrangian. Section III is dedicated to the neutrino mass generation mechanism through the type-III seesaw in our $U(1)_{\rm B-L}$ framework. In Section IV, we explore the DM phenomenology, discussing the relic density calculations. Section V focuses on the direct and indirect detection prospects and constraints from current experiments on the collider phenomenology of our model, with particular emphasis on the production and decay of triplet fermions and the $\rm B-L$ gauge boson. We also discuss the potential for observing disappearing track signatures in future collider experiments. In Section VI, we discuss the gravitational wave signature of the first order phase transition that occurs during the B-L symmetry breaking, providing a complementary probe and finally conclude in Section VII.
II Model
@@
To cancel these remaining anomalies, we introduce two new chiral fermions, $\chi_{L}$ and $\chi_{R}$, with specific $\rm B-L$ charges. The $\rm B-L$ charges of these fermions are calculated to be $(9+\sqrt{57})/6$ and $(-9+\sqrt{57})/6$, respectively [^1] Several explicit models in the literature utilize irrational or fractional $U(1)$ charges in anomaly-free gauge theories. Some illustrative examples include [36, 37, 38, 39, 40, 41, 42]. It is worth noting that such irrational $U(1)$ charges cannot come from a unified gauge theory but remain perfectly consistent in a stand-alone $U(1)$ extension if anomaly cancellation is maintained [36]., ensuring that all three anomaly conditions are simultaneously satisfied. $\chi_{L},\chi_{R}$ constitute a Dirac fermion $\chi$, which serves as the DM candidate in our setup.
@@
$V(H,\Phi_{i},\Phi)$ $=$ $-\mu^{2}_{H}(H^{\dagger}H)+\lambda_{H}(H^{\dagger}H)^{2}+\sum_{i=1}^{3}\big(-\mu^{2}_{\Phi_{i}}(\Phi_{i}^{\dagger}\Phi_{i})+\lambda_{\Phi_{i}}(\Phi_{i}^{\dagger}\Phi_{i})^{2}\big)+(\mu_{1}\Phi_{1}\Phi_{1}\Phi_{2}^{\dagger}+\mu_{2}\Phi_{3}\Phi_{3}\Phi_{1}$ (9)
$+$ $\lambda_{1}\Phi_{3}\Phi_{3}\Phi_{2}\Phi_{1}^{\dagger})+\sum_{i=1}^{3}\lambda_{H\Phi_{i}}(H^{\dagger}H)(\Phi_{i}^{\dagger}\Phi_{i})+\big(-\mu^{2}_{\Phi}(\Phi^{\dagger}\Phi)+\lambda_{\Phi}(\Phi^{\dagger}\Phi)^{2}\big)+\lambda_{H\Phi}(H^{\dagger}H)(\Phi^{\dagger}\Phi)$
$+$ $\mu_{3}\Phi_{1}^{\dagger}\Phi_{3}\Phi+\mu_{4}\Phi_{2}\Phi_{3}\Phi^{\dagger}.$
@@ IV Dark matter
It is worth mentioning here that we used the package micrOMEGAs [43] for computing annihilation cross-section and relic density, after generating the model files using LanHEP [44].
Before delving deeper into the DM phenomenology, it’s crucial to address the constraints on scalar portal couplings arising from Higgs invisible decay searches. In our model, the SM-like Higgs boson $h_{1}$ can potentially decay into pairs of the lighter scalar $h_{2}$ ($h_{1}\to h_{2}h_{2}$) or into DM pairs ($h_{1}\to\chi\bar{\chi}$) through scalar mixing. These decay channels contribute to the Higgs invisible width, which is tightly constrained by experimental searches. The most recent bound on the Higgs invisible decay branching ratio comes from ATLAS collaboration, setting an upper limit of 14.5% at 95% confidence level [45]. This constraint significantly restricts the allowed parameter space for scalar portal couplings and mixing angles as well as DM Yukawa couplings. In our subsequent analysis, we have carefully accounted for this constraint, ensuring that all considered scenarios comply with the current Higgs invisible decay limit. This consideration is crucial as it directly impacts the viable regions of parameter space for DM phenomenology, particularly for DM masses below $M_{h_{1}}/2\simeq 62.5$ GeV.
IV.1 Parameter Space Scan
@@ IV.1.2 Case-II: $M_{h_{2}}\in[1,10^{4}]$ GeV
V Detection Prospects
V.1 Direct Detection
Refer to caption Refer to caption
@@ V.1 Direct Detection
The spin-independent (SI) DM-nucleon cross-section via scalar mediation is given by [46]:
@@
$\mathcal{M}=\mathcal{Z}f_{p}+(A-\mathcal{Z})f_{n},$ (30)
where $A$ is the mass number of the target nucleus, $\mathcal{Z}$ is its atomic number, and $f_{p}$ and $f_{n}$ are the interaction strengths of the DM with protons and neutrons, respectively. These interaction strengths are given by:
@@
The values of $f_{T_{q}}^{p,n}$ and $f_{T_{G}}^{p,n}$ can be found in [47].
The SI DM-nucleon cross-section via $Z_{\rm BL}$ mediation is given by [48]:
$\sigma^{{\rm SI}}_{{\rm vector}}$ $=$ $\frac{\mu_{r}^{2}}{\pi M^{4}_{Z_{\rm BL}}}g_{{}_{\rm BL}}^{2}\bigg(\frac{9+\sqrt{57}}{6}\bigg)^{2}\bigg[\mathcal{Z}(2\tilde{b}_{u}+\tilde{b}_{d})$ (33)
$+$ $(A-\mathcal{Z})(\tilde{b}_{u}+2\tilde{b}_{d})\bigg]^{2},$
@@
Depending on the gauge coupling $g_{\rm BL}$ and the product of the Yukawa coupling $y_{\chi}$ with the mixing angle $\theta_{\phi h}$, either the scalar-mediated process or the vector boson-mediated process dominates, or both become comparable. We compute the DM-nucleon scattering cross-section using micrOMEGAs for parameter points that satisfy the correct relic density constraint and present the results as a function of DM mass for Cases I (IV.1.1) and II (IV.1.2) in Fig. 4, shown in the left and right panels, respectively.
These points are then compared against the most stringent upper limits on the DM-nucleon scattering cross-section from the LUX-ZEPLIN (LZ) direct detection experiment [49], as well as XENONnT [50] and DS-50 [51].
@@
Additionally, we showcase the projected sensitivities of DARWIN [52], SuperCDMS [53], and DS-LM [54], which has the potential to probe a significant portion of the parameter space in future experiments.
V.2 Indirect Detection
Refer to caption Refer to caption
Figure 6: Case-I (left), Case-II (right): DM annihilation cross-section as a function of DM mass for ${\rm DM~DM}\rightarrow\mu^{+}\mu^{-}$ and ${\rm DM~DM}\rightarrow W^{+}W^{-}$ channels. The combined Magic and Fermi-LAT limits [55] are shown with dark cyan and dark red colored dashed lines for $\mu^{+}\mu^{-}$ and $W^{+}W^{-}$ channels, respectively. The projected sensitivity of CTA [56] is shown with the red dotted line for the $W^{+}W^{-}$ channel.
@@ V.2 Indirect Detection
Gamma rays, originating from electromagnetically charged final states, fall within the typical energy range expected for WIMP DM and can be observed using space-based telescopes such as the Fermi Large Area Telescope (Fermi-LAT) or ground-based observatories like MAGIC and HESS. By measuring the gamma-ray flux and incorporating standard astrophysical inputs, one can infer possible DM annihilation channels into various SM final states, such as $\mu^{+}\mu^{-}$, $\tau^{+}\tau^{-}$, and $W^{+}W^{-}$. Since DM particles do not directly couple to photons, the observed gamma rays arise from the decay of these charged particles. Constraints on DM properties can thus be derived by analyzing Fermi-LAT and MAGIC observations of dwarf spheroidal galaxies (dSphs) [55].
Given that DM can annihilate into multiple final states, Fermi-LAT constraints on individual annihilation channels tend to be relatively weak in many scenarios. In Fig. 6, we present the DM annihilation cross-section into $\mu^{+}\mu^{-}$ and $W^{+}W^{-}$ channel as a function of DM mass along with the most stringent constraints from MAGIC+Fermi-LAT. We also show the projected sensitivity of CTA [56] for the $W^{+}W^{-}$ channel with red dotted line. The parameter points shown in this figure correspond to those that satisfy relic density as well as direct detection constraints in both Case I and Case II. Clearly, all points in Fig. 6, which are consistent with relic density requirements as well as direct detection bounds, also remain within the limits set by indirect detection experiments.
V.3 Collider Search

### v3

# Verifiable type-III seesaw and dark matter in a gauged $U(1)_{\rm B-L}$} symmetric model

@@ 1 Introduction
The Standard Model (SM) of particle physics, despite its remarkable success, leaves several fundamental questions unanswered. Two of the most pressing mysteries are the nature of dark matter (DM) and the origin of neutrino masses. Observational evidence from astrophysics and cosmology, including galaxy rotation curves, gravitational lensing, and cosmic microwave background data, strongly suggests that DM constitutes approximately $85\%$ of the Universe’s matter content and $26.8\%$ of its energy density Planck:2018vyg ; ParticleDataGroup:2024cfk . However, the SM lacks a suitable candidate to explain the DM of the universe.
@@
Concurrently, the discovery of neutrino oscillations Super-Kamiokande:1998kpq ; SNO:2001kpb ; DoubleChooz:2011ymz ; DayaBay:2012fng ; RENO:2012mkc has conclusively demonstrated that neutrinos possess non-zero masses, contradicting the SM prediction of massless neutrinos. This observation necessitates an extension of the SM to accommodate neutrino mass generation mechanisms. Various seesaw mechanisms have been proposed to explain the small neutrino masses, including the type I Minkowski:1977sc ; Gell-Mann:1979vob ; Mohapatra:1979ia ; Schechter:1980gr ; yanagida , type II Mohapatra:1980yp ; Lazarides:1980nt ; Wetterich:1981bx ; Schechter:1981cv ; Ma:1998dx , and type III Foot:1988aq seesaw models.
@@
The potential connection between these two fundamental puzzles has sparked considerable interest in the particle physics community. Several beyond Standard Model (BSM) theories have been developed to address both issues within a unified framework. Among these, models based on the gauged $U(1)_{\rm B-L}$ symmetry have garnered significant attention due to their simplicity and potential for experimental verifiability. See for instance  Davidson:1978pm ; Mohapatra:1980qe ; Buchmuller:1991ce ; Sahu:2005fe ; Montero:2007cd ; Ma:2014qra ; Sanchez-Vega:2015qva ; Biswas:2019ygr ; Das:2020uer ; Mahapatra:2020dgk ; Ghosh:2021khk and the references there in.
@@
In this paper, we present a comprehensive study of a $U(1)_{\rm B-L}$ gauged type-III seesaw model that simultaneously addresses the neutrino mass puzzle and provides a DM candidate. We explore the model’s theoretical foundations, analyze its phenomenological consequences, and discuss its implications for current and future experimental searches. The type-III seesaw mechanism Foot:1988aq , which introduces $SU(2)_{L}$ fermion triplets, offers a unique approach to neutrino mass generation. However, when embedded within a gauged $U(1)_{\rm B-L}$ symmetric setup, it introduces new anomalies that require additional particle content for cancellation. Intriguingly, the fermions needed for anomaly cancellation in this setup can serve as viable DM candidates, establishing a natural link between neutrino physics and DM. Specifically, after a careful assignment of $\rm B-L$ charges to the triplet fermions which successfully cancels the $SU(2)_{L}^{2}\times U(1)_{\rm B-L}$ anomalies, two additional chiral fermions are required to cancel the remaining $U(1)_{\rm B-L}^{3}$ and $(grav.)^{2}\times U(1)_{\rm B-L}$ anomalies. These fermions can combine to form a Dirac particle, which gets stabilized by the remnant $Z_{2}$ symmetry after the $U(1)_{\rm B-L}$ breaking and thus becomes a natural DM candidate.
@@
Here, it is worth noting that, as the triplet fermions carry electroweak charge, their production and decay signatures can be prominent at colliders as compared to the type-I seesaw scenario, and by the virtue of their $SU(2)_{L}$ interactions, they can be pair‑produced via electroweak processes with cross sections set by the triplet mass Cai:2017mow ; Jana:2020qzn . In contrast, heavy right-handed neutrinos in type-I seesaw are SM singlets and are only produced via tiny mixing angles, making them essentially invisible at colliders. However, a $\Sigma$ can be copiously produced at the LHC/HL-LHC once kinematically allowed. The charged component $\Sigma^{\pm}$ decays to the neutral $\Sigma^{0}$ plus soft pions (or leptons) with a potentially long lifetime because of the phase space suppression due to the small mass difference arising because of electroweak corrections Cirelli:2005uq . If $\Sigma^{\pm}$ is moderately long-lived, it yields a disappearing track signature in the detector. This striking signal, absent in type-I models, can be searched for at the LHC/HL-LHC. Type-III models also readily produce displaced vertex events, and proposed detectors like MATHUSLA or future lepton–hadron colliders (LHeC, FCC-he) could detect such displaced decays Cai:2017mow ; Jana:2020qzn . Thus, type-III seesaw implementations can be far more testable and constrained than the minimal type-I seesaw.
@@
In addition, the model also predicts unique cosmological signatures in the form of stochastic gravitational waves (GWs). These GWs originate from first-order phase transitions during the breaking of the $U(1)_{\rm B-L}$ symmetry, a process central to the framework’s phenomenology. The resulting GW spectrum can be probed by current and future observatories such as LISA LISA:2017pwj , DECIGO Adelberger:2005bt , and the Einstein Telescope Punturo:2010zz . These experiments provide a complementary avenue to test the model, as the GW signals are sensitive to the dynamics of the phase transition and the underlying particle physics parameters.
@@
The rest of this paper is organized as follows: In Section 2, we provide a detailed description of our model, including the particle content, gauge symmetries, and the Lagrangian. Section 3 is dedicated to the neutrino mass generation mechanism through the type-III seesaw in our $U(1)_{\rm B-L}$ framework. In Section 4, we explore the DM phenomenology, discussing the relic density calculations. Section 5 focuses on the direct and indirect detection prospects and constraints from current experiments on the collider phenomenology of our model, with particular emphasis on the production and decay of triplet fermions and the $\rm B-L$ gauge boson. We also discuss the potential for observing disappearing track signatures in future collider experiments. In Section 6, we discuss the gravitational wave signature of the first order phase transition that occurs during the B-L symmetry breaking, providing a complementary probe and finally conclude in Section 7.
2 Model
@@
To cancel these remaining anomalies, we introduce two new chiral fermions, $\chi_{L}$ and $\chi_{R}$, with specific $\rm B-L$ charges. The $\rm B-L$ charges of these fermions are calculated to be $(9+\sqrt{57})/6$ and $(-9+\sqrt{57})/6$, respectively [^1] Several explicit models in the literature utilize irrational or fractional $U(1)$ charges in anomaly-free gauge theories. Some illustrative examples include Altmannshofer:2019xda ; Kanemura:2014rpa ; Han:2017ars ; Berbig:2022nre ; Nanda:2017bmi ; Gu:2019ird ; Gu:2019gzy . It is worth noting that such irrational $U(1)$ charges cannot come from a unified gauge theory but remain perfectly consistent in a stand-alone $U(1)$ extension if anomaly cancellation is maintained Altmannshofer:2019xda ., ensuring that all three anomaly conditions are simultaneously satisfied. $\chi_{L},\chi_{R}$ constitute a Dirac fermion $\chi$, which serves as the DM candidate in our setup.
@@
$V(H,\Phi_{i},\Phi)$ $=$ $-\mu^{2}_{H}(H^{\dagger}H)+\lambda_{H}(H^{\dagger}H)^{2}+\sum_{i=1}^{3}\big(-\mu^{2}_{\Phi_{i}}(\Phi_{i}^{\dagger}\Phi_{i})$ (9)
$+\lambda_{\Phi_{i}}(\Phi_{i}^{\dagger}\Phi_{i})^{2}\big)+(\mu_{1}\Phi_{1}\Phi_{1}\Phi_{2}^{\dagger}+\mu_{2}\Phi_{3}\Phi_{3}\Phi_{1}$
$+\lambda_{1}\Phi_{3}\Phi_{3}\Phi_{2}\Phi_{1}^{\dagger})+\sum_{i=1}^{3}\lambda_{H\Phi_{i}}(H^{\dagger}H)(\Phi_{i}^{\dagger}\Phi_{i})$
$+\big(-\mu^{2}_{\Phi}(\Phi^{\dagger}\Phi)+\lambda_{\Phi}(\Phi^{\dagger}\Phi)^{2}\big)$
$+\lambda_{H\Phi}(H^{\dagger}H)(\Phi^{\dagger}\Phi)+\mu_{3}\Phi_{1}^{\dagger}\Phi_{3}\Phi$
$+\mu_{4}\Phi_{2}\Phi_{3}\Phi^{\dagger}.$
@@ 4 Dark matter
It is worth mentioning here that we used the package micrOMEGAs Alguero:2023zol for computing annihilation cross-section and relic density, after generating the model files using LanHEP Semenov:2014rea .
Before delving deeper into the DM phenomenology, it’s crucial to address the constraints on scalar portal couplings arising from Higgs invisible decay searches. In our model, the SM-like Higgs boson $h_{1}$ can potentially decay into pairs of the lighter scalar $h_{2}$ ($h_{1}\to h_{2}h_{2}$) or into DM pairs ($h_{1}\to\chi\bar{\chi}$) through scalar mixing. These decay channels contribute to the Higgs invisible width, which is tightly constrained by experimental searches. The most recent bound on the Higgs invisible decay branching ratio comes from ATLAS collaboration, setting an upper limit of 14.5% at 95% confidence level ATLAS:2022yvh . This constraint significantly restricts the allowed parameter space for scalar portal couplings and mixing angles as well as DM Yukawa couplings. In our subsequent analysis, we have carefully accounted for this constraint, ensuring that all considered scenarios comply with the current Higgs invisible decay limit. This consideration is crucial as it directly impacts the viable regions of parameter space for DM phenomenology, particularly for DM masses below $M_{h_{1}}/2\simeq 62.5$ GeV.
4.1 Parameter Space Scan
@@ 4.1.2 Case-II: $M_{h_{2}}\in[1,10^{4}]$ GeV
5 Detection Prospects
5.1 Direct Detection
@@
The spin-independent (SI) DM-nucleon cross-section via scalar mediation is given by Ellis:2008hf :
@@
$\mathcal{M}=Zf_{p}+(A-Z)f_{n},$ (30)
where $A$ is the mass number of the target nucleus, $Z$ is its atomic number, and $f_{p}$ and $f_{n}$ are the interaction strengths of the DM with protons and neutrons, respectively. These interaction strengths are given by:
@@
The values of $f_{T_{q}}^{p,n}$ and $f_{T_{G}}^{p,n}$ can be found in Ellis:2000ds .
The SI DM-nucleon cross-section via $Z_{\rm BL}$ mediation is given by Berlin:2014tja :
$\sigma^{{\rm SI}}_{{\rm vector}}$ $=$ $\frac{\mu_{r}^{2}}{\pi M^{4}_{Z_{\rm BL}}}g_{{}_{\rm BL}}^{2}\bigg(\frac{9+\sqrt{57}}{6}\bigg)^{2}\bigg[Z(2\tilde{b}_{u}+\tilde{b}_{d})$ (33)
$+$ $(A-Z)(\tilde{b}_{u}+2\tilde{b}_{d})\bigg]^{2},$
@@
Depending on the gauge coupling $g_{\rm BL}$ and the product of the Yukawa coupling $y_{\chi}$ with the mixing angle $\theta_{\phi h}$, either the scalar-mediated process or the vector boson-mediated process dominates, or both become comparable. We compute the DM-nucleon scattering cross-section using micrOMEGAs for parameter points that satisfy the correct relic density constraint and present the results as a function of DM mass for Cases I (4.1.1) and II (4.1.2) in Fig. 4, shown in the left and right panels, respectively.
These points are then compared against the most stringent upper limits on the DM-nucleon scattering cross-section from the LUX-ZEPLIN (LZ) direct detection experiment LZ:2024zvo , as well as XENONnT XENON:2023cxc and DS-50 DarkSide-50:2022qzh .
@@
Additionally, we showcase the projected sensitivities of DARWIN DARWIN:2016hyl , SuperCDMS SuperCDMS:2016wui , and DS-LM GlobalArgonDarkMatter:2022ppc , which has the potential to probe a significant portion of the parameter space in future experiments.
5.2 Indirect Detection
Figure 6: Case-I (left), Case-II (right): DM annihilation cross-section as a function of DM mass for ${\rm DM~DM}\rightarrow e^{+}e^{-}$, ${\rm DM~DM}\rightarrow\mu^{+}\mu^{-}$ and ${\rm DM~DM}\rightarrow W^{+}W^{-}$ channels. The combined Magic and Fermi-LAT limits MAGIC:2016xys are shown with dark cyan and dark red colored dashed lines for $\mu^{+}\mu^{-}$ and $W^{+}W^{-}$ channels, respectively. The projected sensitivity of CTA CTA:2020qlo is shown with the red dotted line for the $W^{+}W^{-}$ channel. CMB constraints Slatyer:2015jla are shown with blue dashed and dark cyan dashed lines for $e^{+}e^{-}$ and $\mu^{+}\mu^{-}$ channels, respectively.
@@ 5.2 Indirect Detection
Gamma rays, originating from electromagnetically charged final states, fall within the typical energy range expected for WIMP DM and can be observed using space-based telescopes such as the Fermi Large Area Telescope (Fermi-LAT) or ground-based observatories like MAGIC and HESS. By measuring the gamma-ray flux and incorporating standard astrophysical inputs, one can infer possible DM annihilation channels into various SM final states, such as $\mu^{+}\mu^{-}$, $\tau^{+}\tau^{-}$, and $W^{+}W^{-}$. Since DM particles do not directly couple to photons, the observed gamma rays arise from the decay of these charged particles. Constraints on DM properties can thus be derived by analyzing Fermi-LAT and MAGIC observations of dwarf spheroidal galaxies (dSphs) MAGIC:2016xys .
Given that DM can annihilate into multiple final states, Fermi-LAT constraints on individual annihilation channels tend to be relatively weak in many scenarios. In Fig. 6, we present the DM annihilation cross-section into $\mu^{+}\mu^{-}$ and $W^{+}W^{-}$ channel as a function of DM mass along with the most stringent constraints from MAGIC+Fermi-LAT. We also show the projected sensitivity of CTA CTA:2020qlo for the $W^{+}W^{-}$ channel with red dotted line. The parameter points shown in this figure correspond to those that satisfy relic density as well as direct detection constraints in both Case I and Case II. Clearly, all points in Fig. 6, which are consistent with relic density requirements as well as direct detection bounds, also remain within the limits set by indirect detection experiments.
In addition to indirect search constraints, it is also important to consider the bounds arising from CMB measurements, which are sensitive to energy injection from DM annihilation around the time of recombination. These constraints are usually expressed through the quantity
$p_{\rm ann}=f_{\rm eff}(z)\,\frac{\langle\sigma v\rangle}{M_{\rm DM}}$
where $f_{\rm eff}(z)$ denotes the redshift-dependent efficiency factor corresponding to a given annihilation channel. The Planck analysis places the upper bound
$p_{\rm ann}\lesssim 4.1\times 10^{-28}~{\rm cm^{3}\,s^{-1}\,GeV^{-1}}$
at $95\%$ C.L. Slatyer:2015jla . In Fig. 6, we show the corresponding CMB limits for the $e^{+}e^{-}$ and $\mu^{+}\mu^{-}$ channels (blue and dark-cyan dashed lines). Clearly, all points satisfying the correct relic density and direct detection constraints have annihilation cross sections into $e^{+}e^{-}$ that lie well below the CMB exclusion band, and hence are consistent with Planck bounds. We also note that at lower DM masses, where both the CMB and indirect detection limits are particularly stringent, often approaching or excluding the canonical thermal cross-section, our scenario still remains viable. This is because the relic density in this mass range is dominantly determined by the DM annihilation into $h_{2}h_{2}$ via a $t$-channel process, while the branching to other channels is small. As DM annihilation to $h_{2}h_{2}$ is a $p$-wave-suppressed process, the thermally averaged annihilation cross-section consequently decreases with temperature, making the late-time annihilation rate negligibly small and allowing the model to evade both CMB and indirect detection constraints.
5.3 Collider Search